Find the angle between the lines whose direction cosines satisfy the equations l+m+n=0 and 2nl+2lm-mn=0.
1
Expert's answer
2019-09-12T05:10:32-0400

#Task Q94214

"l+m+n=0"

"m=-(l+n)"

"2lm+2ln-mn=0"

"2ln+m(2l-n)=0"

"2ln-(l+n)(2l-n)=0"

"2ln-(2l^2-nl=2ln-n^2 )=0"

"-2l^2+nl+n^2=0"

"2l^2-nl-n^2=0"

"2l^2-2nl+nl-n^2=0"

"2l(l-n)+n(l-n)=0"

"(l-n)(2l+n)=0"

"l=n"

"l=-n\/2"

"m=-(n+n)=-2n"

"n= -m\/2"

"m=-(-n\/2+n)"

"m=-n\/2"

"l=-m\/2=n\/1"

"l\/1=m\/1=-n\/2"

"l:m:n=1:-2:1"

"l:m:n=1:1:-2"

"cos \theta=\frac{1*1-2*1-2*1}{\sqrt{1^2+2^2+1^2}\sqrt{1^2+1^2+2^2}}=-\frac{1}{2}"

"\u03b8=120^0"